OpenStudy (b.r):

1. What are the next three terms in the sequence? –3, 6, 15, 24, … 35, 46, 57 34, 44, 54 33, 44, 56 33, 42, 51

OpenStudy (domebotnos):

33 42 51

OpenStudy (anonymous):

these are in A.P. first term a= -3 c.d. d=6-(-3)=6+3=9=15-6 \[t _{n}=a+(n-1)d\]